OpenStudy (anonymous):

Ticket to a community theater cost $25 for adults and $5 for children. At a recent performance a total of 305 tickets were sold for $6025. Determine how many tickets of each kind were sold.

OpenStudy (anonymous):

Let the number of adult tickets sold be x. So, number of child tickets sold = 305 - x. 25x + 5(305 - x) = 6025 25x + 1525 - 5x = 6025 20x = 6025 - 1525 = 4500 x = 4500/20 = 225 305 - x = 305 - 225 = 80

OpenStudy (anonymous):

thank you pratu042

OpenStudy (anonymous):

043